OpenStudy (anonymous):

graph the inequality on a plane 2x+3y<=to6

OpenStudy (anonymous):

What would 2x + 3y = 6 look like?

OpenStudy (anonymous):

(I think they're only dealing with \(R^2\) here)

OpenStudy (anonymous):

Let's forget about the inequality for a while

OpenStudy (anonymous):

2x+3y=6 3y=-2x+6 y=-2/3x +2

OpenStudy (anonymous):

thanks my pal

OpenStudy (anonymous):

y<=-2/3x+2